OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined. q^3-4q/q^2-49

OpenStudy (anonymous):

First off please write your expression in a more simplistic manner, because this form is hard to comprehend. Assuming you were going for (q^3-4q)/(q^2-49): = (q^2-49) = (q-7)*(q+7) meaning you have undefined portions in the function where your denominator is equal to 0. In this case the denominator is equal to zero when q=\[\pm7\]. those are the only ones that i see since cubics are smooth functions and have no undefined portions.